Admission to a national park is so much per van. A group of friends in a van shared the admission cost by paying $0.50 apiece. If 2 more friends had joined them, everyone would have paid on $0.40 apiece. How much does it cost a van to enter the park?
Answer by checkley71(8403)

X(.50)=(X+2).40 OR .50X=.40X+.80 OR (.50-.40)X=.80 OR .1X=.80 OR X=.80/.1 OR
X=8.00 IS THE NUMBER OF PEOPLE PAYING $.50 THUS THE TOTAL PRICE FOR THE VAN IS 8*.50 OR $4.00 THUS (8+2).40=$4.00 IF 10 PEOPLE EACH PAY $.40